# The Landscape of GPU-Centric Communication

### v2

@@ 5.1. Moving Away from CPU
There are several reasons why we believe moving away from CPU-controlled execution shows promise. First, it addresses the issue of CPU-induced latency barriers that are caused by kernel launch and memory copy overheads. These barriers become more significant in strong scaling scenarios as the number of GPUs increases and computation per GPU decreases. In latency-bound settings, traditional CPU-controlled implementations are not able to overlap communication with computation as the latencies to initiate the operations take longer than the operations themselves effectively serializing communication and computation. On the other hand, CPU-free execution can still achieve adequate levels of overlap even when latencies dominate (Wagner, 2020; Ismayilov et al., 2023). Second, the parallelism offered by within-kernel communication is well-suited for persistent kernels to take advantage of. With the use of efficient communication and synchronization APIs, this execution model can achieve higher bandwidths and lower latencies than CPU-controlled implementations. Additionally, CPU-free execution, by virtue of inlining communication with computation, is well-suited for applications with fine-grained communication (Chen et al., 2022a). Third, ROC_SHMEM allows for the first time to fully migrate application execution to the device with zero reliance on CPU helper threads. NVSHMEM with its IBGDA transport can also migrate a significant amount of execution to GPU but must still rely on the CPU for functional correctness.
@@
However, there are several challenges facing this model of execution. One major challenge is that persistent kernels can result in reduced occupancy potentially bottlenecking computation. As of today, if global device or multi-GPU barriers are required, persistent kernels must be launched in a cooperative manner. This means that only as many threads can be launched as can run concurrently at the same time making hardware oversubscription impossible. As a result, workload decomposition and scheduling, which were previously handled by the hardware scheduler, now need to be manually done by the programmer. This manual approach is unlikely to be as efficient as hardware-based scheduling, and compute-intensive applications are likely to suffer. Furthermore, long-running persistent kernels will consume more registers and may use shared memory as well limiting occupancy even further. Nevertheless, we see a two-fold solution to this problem. First, there is a large amount of high bandwidth shared memory available across application execution, which can potentially nullify the performance hit caused by reduced occupancy. Second, we predict that as GPU vendors strive more and more for greater GPU autonomy, they will introduce APIs that allow for combining persistent kernels with hardware oversubscription. The manual decomposition can also be handled by an optimized compiler / runtime system.
@@ 5.2. UCX as a Potential Pathway for GPU-Awareness
Unified Communication X (UCX) is an open-source communication framework that abstracts over several network APIs, programming models, protocols, and implementations. The idea is to provide a set of high-level primitives while hiding the low-level implementation details behind the UCX runtime. The relevance of UCX for GPU-centric communication is that its tagged and stream APIs can be used to implement a GPU-centric communication layer for both ROCm and CUDA. The actual GPU-centric communication then uses the corresponding native libraries (Shamis et al., 2015; The Unified Communication X Library, [n. d.]).
Using UCX for realizing GPU-centric communication is a recent direction in the literature that has started to take hold. Perhaps the most relevant example is that of ROCm-awareness for MPI implementations. Much early work on GPU-aware MPI was done for NVIDIA GPUs using native CUDA libraries and then integrated directly into MPI runtimes. Perhaps reticent to replicate the same work to make their runtimes ROCm-aware, most MPI implementations provide ROCm-awareness only through UCX. OpenMPI additionally provides CUDA support through UCX, besides its native integration. In non-MPI work, Choi et al. extend the UCX layer in Charm++ to provide GPU-aware communication for several programming models in the Charm++ ecosystem (Choi et al., 2021a, 2022).
We predict that more runtimes will gravitate toward UCX to add support for GPU-centric communication. Using UCX APIs frees the programmers from relying on native vendor-specific APIs and allows adding GPU-aware communication for both ROCm and CUDA. However, there are some caveats. One is that it needs to be clarified if GPU-aware UCX can overlap communication with computation. The second caveat is that the increased generality and convenience may trade off performance. A performance comparison between GPU-awareness through native APIs and that provided through UCX would be helpful. In one work in this direction, Khorassani et al. provide a native ROCm-aware runtime for MVAPICH2, which outperforms OpenMPI with UCX on a cluster of AMD GPUs (Shafie Khorassani et al., 2021). However, the performance difference may be due to differences in the MPI implementations, not UCX. Another current limitation of UCX is that it is always on the host, and it does not use any communication kernel, but relies on the traditional cudamemcpy family of functions in addition to zero copy RDMA.
@@ 5.3. CPU-Free Networking
In the earliest work, GGAS (Oden and Fröning, 2013) proposes changes to network devices to implement a unified global address space that allows moving the control path entirely to the GPU. This is accomplished by using a persistent kernel that contains the computation, communication, and synchronization all on the device-side. While the work was the first of its kind and showed performance improvements compared to a CUDA-Aware MPI baseline, the experiments were conducted on two GPU nodes with one GPU each, while the proposed hardware changes were emulated on an FPGA. Follow-up work showed that GGAS, by virtue of eliminating CPU involvement in the control path, can achieve further performance improvements and reduce energy usage compared to CPU-controlled baselines (Oden et al., 2014b; Klenk et al., 2014b).
@@
Several more works on CPU-free networking followed. Oden et al. (Oden et al., 2014a) use GPUDirect RDMA to allow GPUs to directly interface with Infiniband network devices without the involvement of the host CPU. They do this by mapping the entire Infiniband context to the device-side and using the GPU to generate and send work requests to the HCA. However, because of slow single-thread work request generation performance on the GPU, the proposed changes deteriorated performance compared to CPU-controlled baselines. Follow-up work ameliorated these performance limitations and showed much more promising results (Klenk et al., 2014a, 2015). Another work combines the proposed GPU-side Infiniband Verbs with CUDA Dynamic Parallelism to optimize the bottleneck of intra-kernel synchronization (Oden et al., 2014c). GPUrdma also implements Infiniband on the GPU and proposes a GPU-side library for direct communication from within persistent GPU kernels with zero CPU involvement. The proposed design outperforms a CPU-controlled baseline on a series of microbenchmarks but runs into correctness issues resulting from the use of persistent kernels and GPU-NIC interaction (Daoud et al., 2016).
@@
Silberstein et al. implement GPUNet, which provides GPU-side socket abstractions and networking primitives (Silberstein et al., 2016). GPUNet allows invoking the communication on the GPU but does not fully migrate the control path to the device; instead, it relies on CPU helper threads to perform the actual communication. A similar approach is adopted by dCUDA, which provides device-side APIs with MPI semantics but translates them to standard MPI calls performed by CPU helper threads (Gysi et al., 2016). LeBeane et al. categorized GPU networking methods discussing at length their deficiencies. In response, they propose GPU-TN, a NIC hardware mechanism that allows the CPU to create and register messages with the NIC and the GPU to trigger them from a running persistent kernel (LeBeane et al., 2017). Another work, ComP-Net, uses embedded GPU microprocessors to offload helper threads from the CPU to the GPU (LeBeane et al., 2018). While both GPU-TN and ComP-Net show promising performance, they require hardware changes to the NIC and the GPU and, thus, rely on simulation to obtain results.
@@ 5.4. Broader GPU Autonomy
The recent proliferation of GPU-centric communication represents the general trend toward broader GPU autonomy. Several works, early and recent, have tried to hand the GPU the reins of domains that have traditionally been the purview of the CPU. In an early work, Stuart et al. propose methods that allow the GPU to issue callbacks to the CPU (Stuart et al., 2010). Silberstein et al. implement GPUfs, which allows the GPU to request files on the host CPU directly from inside a GPU kernel (Silberstein et al., 2014). Veselý et al. implement support for invoking POSIX system calls from inside GPU kernels through changes to the Linux kernel (Veselý et al., 2018). NVIDIA’s GPUDirect Storage provides a direct data path between GPUs and storage but still relies on the CPU to orchestrate execution (Thompson and C.J., 2012). In a recent work by NVIDIA, Qureshi et al. present BaM, the first approach that allows GPUs to directly access storage without any CPU involvement. Experimental results show that BaM outperforms GPUDirect Storage on several workloads (Qureshi et al., 2023).
These and other works show a clear trend toward general GPU autonomy. In line with this, we expect further optimizations to GPU-centric communication. Several recent mechanisms are promising, as pointed out by Punniyamurthy et al. (Punniyamurthy et al., 2023a). First, the recent thread block(TB) cluster abstraction could benefit device-side communication-computation overlap and inter-TB synchronization. Second, AMD’s recent cache flush instructions allow flushing the cache before initiating network communication, meaning communication no longer needs to be allocated as uncacheable. Third, recent hardware trends like fatter GPU nodes and tight GPU-NIC integration are also promising (Evans et al., 2022; AMD, 2021). For example, the most recent iteration of NVSwitch directly connects 256 Grace Hopper Superchips enabling direct P2P all-to-all communication at an unprecedented scale.
@@ 5.5. Collective Algorithms Design
Multi-GPU systems present unprecedented challenges in the design of collective operations algorithms. First, GPUs within the same node are often arranged in complex and non-traditional topologies, making existing collective algorithms less effective. Second, there is significant heterogeneity in the bandwidth between GPU pairs, both within a node and between nodes. For instance, there can be up to a 4x difference in bandwidth among different GPU pairs on the same node (CSC, 2024), and up to a 10x gap between intra-node and inter-node bandwidth. Third, these topological and connectivity characteristics frequently change between GPU generations. These factors complicate the design of collective algorithms, potentially leading to underutilization of available bandwidth and poor performance of collective operations.
Some approaches use linear programming formulations to find the optimal collective algorithm based on a network specification and the size of the collective (Shah et al., 2022; Cai et al., 2021; Wang et al., 2019). However, this involves solving an NP-hard problem that scales exponentially. For example, finding a solution for 128 nodes can take up to 11 hours (Shah et al., 2022), and a new solution might be needed if the number of nodes or the size of the collective changes. This complexity makes generating collective algorithms for large systems challenging, if not impossible. To simplify the implementation of multi-GPU collective operations, MSCCL (Cowan et al., 2023) provides a high-level language for expressing collective operations, which is then compiled to NCCL code.
5.6. Debugging and Profiling Support
@@
Palwisha et al. present ComScribe (Akhtar et al., 2021), a tool that allows monitoring NCCL collective and P2P communication. However, it is limited to a single node and relies on the deprecated nvprof tool. New efforts in this area have been made by Snoopie (Issa et al., 2024), which focuses on profiling and visualizing GPU-centric communication. Snoopie is capable of attributing communication to source code lines and objects involved in communication, offering different levels of granularity. This allows for both a coarse-grained overview of the system and a detailed view of a specific object or device in terms of data movement.
Another class of tools that can help with debugging GPU communication are race detectors. Given that many of the aforementioned communication libraries use a partitioned global address space model compared with message passing, there is a high likelihood of introducing race hazards. Race detector tools are vital to navigate the complexities of shared data in multi-GPU programming. Despite their importance, none of the existing tools are capable of detecting race hazards in multi-GPU programming. Compute Sanitizer’s Racecheck tool (NVIDIA, [n. d.]) is limited to on-chip shared memory, supporting the detection of race hazards only within a single GPU context. HiRace (Jacobson et al., 2024) improves on that by supporting global memory and can detect many types of race hazards that Compute Sanitizer’s Racecheck cannot, but it is also limited to a single GPU context.

### v3

@@ 5.1. Moving Away from CPU
There are several reasons why we believe moving away from CPU-controlled execution shows promise. First, it addresses the issue of CPU-induced latency barriers that are caused by kernel launch and memory copy overheads. These barriers become more significant in strong scaling scenarios as the number of GPUs increases and computation per GPU decreases. In latency-bound settings, traditional CPU-controlled implementations are not able to overlap communication with computation as the latencies to initiate the operations take longer than the operations themselves effectively serializing communication and computation. On the other hand, CPU-free execution can still achieve adequate levels of overlap even when latencies dominate (Talk-NVSHMEM-Overcoming-Latency-Barriers; Multi-GPU-Communication-Schemes-for-Iterative-Solvers-When-CPUs-Are-Not-in-Charge). Second, the parallelism offered by within-kernel communication is well-suited for persistent kernels to take advantage of. With the use of efficient communication and synchronization APIs, this execution model can achieve higher bandwidths and lower latencies than CPU-controlled implementations. Additionally, CPU-free execution, by virtue of inlining communication with computation, is well-suited for applications with fine-grained communication (Atos-Vyse-Scalable-Irregular-Parallelism-with-GPUs-Getting-CPUs-Out-of-the-Way). Third, ROC_SHMEM allows for the first time to fully migrate application execution to the device with zero reliance on CPU helper threads. NVSHMEM with its IBGDA transport can also migrate a significant amount of execution to GPU but must still rely on the CPU for functional correctness.
@@
However, there are several challenges facing this model of execution. One major challenge is that persistent kernels can result in reduced occupancy potentially bottlenecking computation. As of today, if global device or multi-GPU barriers are required, persistent kernels must be launched in a cooperative manner. This means that only as many threads can be launched as can run concurrently at the same time making hardware oversubscription impossible. As a result, workload decomposition and scheduling, which were previously handled by the hardware scheduler, now need to be manually done by the programmer. This manual approach is unlikely to be as efficient as hardware-based scheduling, and compute-intensive applications are likely to suffer. Furthermore, long-running persistent kernels will consume more registers and may use shared memory as well limiting occupancy even further. Nevertheless, we see a number of solution to this problem. First, there is a large amount of high bandwidth shared memory available across application execution, which can potentially nullify the performance hit caused by reduced occupancy. Second, we predict that as GPU vendors strive more and more for greater GPU autonomy, they will introduce APIs that allow for combining persistent kernels with hardware oversubscription. The manual decomposition can also be handled by an optimized compiler / runtime system. Alternatively, collective operations can be offloaded to the network components (parallelkittens), freeing the resources in the GPU to compute.
@@ 5.2. UCX as a Potential Pathway for GPU-Awareness
Unified Communication X (UCX) is an open-source communication framework that abstracts over several network APIs, programming models, protocols, and implementations. The idea is to provide a set of high-level primitives while hiding the low-level implementation details behind the UCX runtime. UCX is organized into three main components: UCP (UC-Protocol), which provides high-level communication primitives such as messaging and remote memory access; UCT (UC Transport), which offers low-level access to network hardware; and UCS (UC Services), which supplies common utilities for memory management and threading.
At runtime, a UCP layer dynamically selects the appropriate UCT transport based on the pointer type and system topology. For example, when a CUDA pointer is passed to an inter-node UCP communication, the rc_mlx5 transport may be chosen, whereas for intra-node communication, the cuda_ipc transport is typically preferred. As a result, UCP can accept device pointers as data payloads and perform transfers using the most suitable mechanism, making UCX inherently GPU-aware. This GPU-awareness of UCX allows higher-level communication libraries such as OpenMPI and MPICH to handle device pointers directly in their communication primitives. For example in CUDA-aware MPI, UCX eliminates explicit host memory copies by enabling GPU memory pointers to be passed directly to MPI functions (UCX-an-open-source-framework-for-HPC-network-APIs-and-beyond; openucx-website).
Using UCX for realizing GPU-centric communication is a recent direction in the literature that has started to take hold. Perhaps the most relevant example is that of ROCm-awareness for MPI implementations. Much early work on GPU-aware MPI was done for NVIDIA GPUs using native CUDA libraries and then integrated directly into MPI runtimes. Perhaps reticent to replicate the same work to make their runtimes ROCm-aware, most MPI implementations provide ROCm-awareness only through UCX. OpenMPI and MPICH additionally provide CUDA support through UCX, besides their native integration. In non-MPI work, Choi et al. extend the UCX layer in Charm++ to provide GPU-aware communication for several programming models in the Charm++ ecosystem (Charm-GPU-Aware-1-GPU-Aware-Communication-with-UCX-in-Parallel-Programming-Models; Charm-GPU-Aware-2-Accelerating-communication-for-parallel-programming-models-on-GPU-systems).
We predict that more runtimes will gravitate toward UCX to add support for GPU-centric communication. Using UCX APIs frees the programmers from relying on native vendor-specific APIs and allows adding GPU-aware communication for both ROCm and CUDA. A performance comparison between GPU-awareness through native APIs and that provided through UCX would be helpful. In one work in this direction, Khorassani et al. provide a native ROCm-aware runtime for MVAPICH2, which outperforms OpenMPI with UCX on a cluster of AMD GPUs (Designing-a-ROCm-Aware-MPI-Library-for-AMD-GPUs-Early-Experiences). However, the performance difference may be due to differences in the MPI implementations, not UCX. A current limitation of UCX is that it is always on the host, and it does not use any communication kernel, but relies on the traditional cudamemcpy family of functions in addition to zero copy RDMA.
@@ 5.3. CPU-Free Networking
In the earliest work, GGAS (GGAS-Global-GPU-Adress-Spaces-For-Efficient-Communication-in-Heterogeneous-Clusters) proposes changes to network devices to implement a unified global address space that allows moving the control path entirely to the GPU. This is accomplished by using a persistent kernel that contains the computation, communication, and synchronization all on the device-side. While the work was the first of its kind and showed performance improvements compared to a CUDA-Aware MPI baseline, the experiments were conducted on two GPU nodes with one GPU each, while the proposed hardware changes were emulated on an FPGA. Follow-up work showed that GGAS, by virtue of eliminating CPU involvement in the control path, can achieve further performance improvements and reduce energy usage compared to CPU-controlled baselines (GGAS-Energy-Efficient-Collective-Reduce-and-Allreduce-Operations-on-Distributed-GPUs; GGAS-GPU-Centric-Communication-for-Improved-Efficiency).
@@
Several more works on CPU-free networking followed. Oden et al. (Infiniband-Verbs-on-GPU-A-Case-Study-of-Controlling-an-Infiniband-Network-Device-from-the-GPU) use GPUDirect RDMA to allow GPUs to directly interface with Infiniband network devices without the involvement of the host CPU. They do this by mapping the entire Infiniband context to the device-side and using the GPU to generate and send work requests to the HCA. However, because of slow single-thread work request generation performance on the GPU, the proposed changes deteriorated performance compared to CPU-controlled baselines. Follow-up work ameliorated these performance limitations and showed much more promising results (Analyzing-Put-Get-APIs-for-Thread-Collaborative-Processors; Analyzing-Communication-Models-for-Distributed-Thread-Collaborative-Processors-in-Terms-of-Energy-and-Time). Another work combines the proposed GPU-side Infiniband Verbs with CUDA Dynamic Parallelism to optimize the bottleneck of intra-kernel synchronization (Energy-Efficient-Stencil-Computations-on-Distributed-GPUs-Using-Dynamic-Parallelism-and-GPU-Controlled-Communication). GPUrdma also implements Infiniband on the GPU and proposes a GPU-side library for direct communication from within persistent GPU kernels with zero CPU involvement. The proposed design outperforms a CPU-controlled baseline on a series of microbenchmarks but runs into correctness issues resulting from the use of persistent kernels and GPU-NIC interaction (GPUrdma-GPU-Side-Library-for-High-Performance-Networking-from-GPU-Kenrels).
@@
Silberstein et al. implement GPUNet, which provides GPU-side socket abstractions and networking primitives (GPUNet-Networking-Abstractions-for-GPU-Programs). GPUNet allows invoking the communication on the GPU but does not fully migrate the control path to the device; instead, it relies on CPU helper threads to perform the actual communication. A similar approach is adopted by dCUDA, which provides device-side APIs with MPI semantics but translates them to standard MPI calls performed by CPU helper threads (dCUDA-Hardware-Supported-Overlap-of-Computation-and-Communication). LeBeane et al. categorized GPU networking methods discussing at length their deficiencies. In response, they propose GPU-TN, a NIC hardware mechanism that allows the CPU to create and register messages with the NIC and the GPU to trigger them from a running persistent kernel (GPU-TN-GPU-Triggered-Networking-for-Intra-Kernel-Communication). Another work, ComP-Net, uses embedded GPU microprocessors to offload helper threads from the CPU to the GPU (ComP-Net-Command-Processor-Networking-for-Efficient-Intra-Kernel-Communication-on-GPUs). While both GPU-TN and ComP-Net show promising performance, they require hardware changes to the NIC and the GPU and, thus, rely on simulation to obtain results.
@@ 5.4. Broader GPU Autonomy
The recent proliferation of GPU-centric communication represents the general trend toward broader GPU autonomy. Several works, early and recent, have tried to hand the GPU the reins of domains that have traditionally been the purview of the CPU. In an early work, Stuart et al. propose methods that allow the GPU to issue callbacks to the CPU (GPU-to-CPU-Callbacks). Silberstein et al. implement GPUfs, which allows the GPU to request files on the host CPU directly from inside a GPU kernel (GPUfs-Integrating-a-File-System-with-GPUs). Veselý et al. implement support for invoking POSIX system calls from inside GPU kernels through changes to the Linux kernel (Generic-System-Calls-for-GPUs). NVIDIA’s GPUDirect Storage provides a direct data path between GPUs and storage but still relies on the CPU to orchestrate execution (CUDA-Blog-GPUDirect-Storage-A-Direct-Path-Between-Storage-and-GPU-Memory).
SmartIO enables low-cost, efficient I/O disaggregation across PCIe-connected hosts by allowing remote machines to borrow and directly access devices such as NVMe, GPUs, and NICs as if they were local. Building on this idea, Qureshi et al. at NVIDIA present BaM allowing GPUs to directly access storage without any CPU involvement. Experimental results show that BaM outperforms GPUDirect Storage on several workloads (GPU-Initiated-On-Demand-High-Throughput-Storage-Access-in-the-BaM-System-Architecture). Turimbetov et al. showed CPU-free device-side task graph execution on multiple GPUs, eliminating kernel launch and scheduling overheads from the host-side (mustard). Lastly, Baydamirli et al. (Autonomous-Execution-for-Multi-GPU-Systems-Compiler-Support) presented a compiler support for autonomous execution for multi-GPU systems, enabling GPU-initiated communication within Python code using NVSHMEM.
These and other works show a clear trend toward general GPU autonomy. In line with this, we expect further optimizations to GPU-centric communication. Several recent mechanisms are promising, as pointed out by Punniyamurthy et al. (GPU-initiated-Fine-grained-Overlap-of-Collective-Communication-with-Computation). First, the recent thread block (TB) cluster abstraction could benefit device-side communication-computation overlap and inter-TB synchronization. Second, AMD’s recent cache flush instructions allow flushing the cache before initiating network communication, meaning communication no longer needs to be allocated as uncacheable. Third, recent hardware trends like fatter GPU nodes and tight GPU-NIC integration are also promising (CUDA-Blog-NVIDIA-Grace-Hopper-Superchip-Architecture-In-Depth; AMD-CDNA2-Architecture-Whitepaper). For example, the most recent iteration of NVSwitch directly connects 256 Grace Hopper Superchips enabling direct P2P all-to-all communication at an unprecedented scale.
While these advances push the system toward broader GPU autonomy, the CPU plays a complementary role to provide system-level responsibilities. Debugging, profiling, monitoring, and software development still fundamentally rely on the CPU’s broader visibility into the system, its mature tooling ecosystem, and its tight integration with operating system services.
@@ 5.5. Collective Algorithms Design
Multi-GPU systems present unprecedented challenges in the design of collective operations algorithms. First, GPUs within the same node are often arranged in complex and non-traditional topologies, making existing collective algorithms less effective. Second, there is significant heterogeneity in the bandwidth between GPU pairs, both within a node and between nodes. For instance, there can be up to a 4x difference in bandwidth among different GPU pairs on the same node (lumig), and up to a 10x gap between intra-node and inter-node bandwidth. Third, these topological and connectivity characteristics frequently change between GPU generations. These factors complicate the design of collective algorithms, potentially leading to underutilization of available bandwidth and poor performance of collective operations.
Some approaches use linear programming formulations to find the optimal collective algorithm based on a network specification and the size of the collective (shah2022taccl; 10.1145/3437801.3441620; Blink-Fast-and-Generic-Collectives-for-Distributed-ML). However, this involves solving an NP-hard problem that scales exponentially. For example, finding a solution for 128 nodes can take up to 11 hours (shah2022taccl), and a new solution might be needed if the number of nodes or the size of the collective changes. This complexity makes generating collective algorithms for large systems challenging, if not impossible. To simplify the implementation of multi-GPU collective operations, MSCCL (10.1145/3575693.3575724) provides a high-level language for expressing collective operations, which is then compiled to NCCL code.
5.6. Debugging, Profiling, Benchmarking Support
@@
Palwisha et al. present ComScribe (ComScribe-Identifying-Intra-Node-GPU-Communication), a tool that allows monitoring NCCL collective and P2P communication. However, it is limited to a single node and relies on the deprecated nvprof tool. New efforts in this area have been made by Snoopie (Snoopie), which focuses on profiling and visualizing GPU-centric communication. Snoopie is capable of attributing communication to source code lines and objects involved in communication, offering different levels of granularity. This allows for both a coarse-grained overview of the system and a detailed view of a specific object or device in terms of data movement.
Another class of tools that can help with debugging GPU communication are race detectors. Given that many of the aforementioned communication libraries use a partitioned global address space model compared with message passing, there is a high likelihood of introducing race hazards. Race detector tools are vital to navigate the complexities of shared data in multi-GPU programming. Despite their importance, none of the existing tools are capable of detecting race hazards in multi-GPU programming. Compute Sanitizer’s Racecheck tool (Racecheck) is limited to on-chip shared memory, supporting the detection of race hazards only within a single GPU context. HiRace (HiRace) improves on that by supporting global memory and can detect many types of race hazards that Compute Sanitizer’s Racecheck cannot, but it is also limited to a single GPU context.
Several benchmarking tools have been proposed for measuring the performance of GPU-GPU communication. Standard suites such as the *ccl-tests for NCCL and RCCL are commonly used to assess the performance of both collectives and point-to-point communication (nccl-tests; rccl-tests). Similarly, the OSU benchmark suite (OMB) provides extensive support for GPU-aware MPI environments (10.1007/978-3-642-33518-1_16). NVSHMEM also provides specific micro-benchmarks to evaluate device-initiated put and get latency. However, these tools typically report performance measurements aggregated across multiple iterations, which masks critical network performance variability and transient noise (sc2019). Furthermore, they often lack baselines for explicit peer-to-peer copies. To address these limitations, the Blink-GPU benchmark was proposed (gpugpuinterconnect; blink-gpu), enabling the capture of per-iteration timing to expose performance anomalies that aggregated metrics miss.
@@
5.7. Compression-Accelerated GPU Communication
The high cost of moving large data volumes across the network, particularly for inter-node GPU communication, has motivated the development of techniques for compressing data prior to transmission. While this approach introduces computational overhead for compression and decompression, this cost is frequently amortized by the substantial reduction in communication time. To maximize the data reduction (i.e., achieve a high compression ratio), lossy compression is typically employed. Specifically, error-bounded lossy compression (10.1145/3733104) allows for high compression ratios while guaranteeing that data distortion remains within a user-defined accuracy threshold. This technique has been integrated into collective communication libraries for both CPU and GPU architectures (10793125; 10.1145/3650200.3656636; 10.1145/3721145.3733642; 10.1007/978-3-031-07312-0_1; coccl), demonstrating significant speedups over conventional libraries like MPI and NCCL.
While effective at reducing data volume, these first-generation solutions were fundamentally limited by the Decompression-Operation-Compression (DOC) workflow. This workflow mandates that data be fully decompressed before any computation (such as a reduction in an Allreduce operation) can be performed, only to be immediately recompressed for subsequent steps, incurring significant processing overhead. This limitation has motivated the development of homomorphic compression, which enables computation to be performed directly on compressed data. Recent approaches have successfully implemented this paradigm for both CPU (10793125) and GPU (10.1145/3721145.3733642) architectures. By allowing GPUs to compute and communicate entirely in the compressed domain, this approach represents the most advanced solution, maximizing throughput by addressing both the data volume bottleneck and the internal DOC processing overhead.
5.8. Level of Maturity and User-Level Perspective
It is crucial to clarify the role and maturity of the technologies that underpin modern accelerator communication, such as NVIDIA’s GPUDirect, GPUDirect RDMA, AMD’s ROCm peer-to-peer capabilities, and GDRCopy. These components are not high-level, user-facing APIs, nor are they hacks. Rather, they are foundational, vendor-supported hardware and driver-level functionalities. They represent the official, low-level architecture integrated directly into the accelerator’s core runtime environment and the system’s kernel drivers.
The maturity and reliability of these solutions are demonstrated by their universal adoption as the essential building blocks for the entire high-performance, open-source communication ecosystem. High-level libraries, including all major MPI implementations (e.g., Open MPI, MVAPICH, Intel MPI), low-level communication middleware (e.g., UCX), and vendors’s own *CCL, are built on top of these APIs. These libraries explicitly use those lower-level functionalities to enable their GPU-aware communication paths. Therefore, the benefits of these technologies (namely, the elimination of the CPU as a bottleneck and the ability to achieve true, zero-copy, low-latency communication by bypassing host memory) are the well-documented, stable, fundamental mechanisms that enable all modern multi-node, multi-GPU supercomputing.